OpenStudy (anonymous):

For y = x2 + 6x - 16, Determine if the parabola opens up or down. State if the vertex will be a maximum or minimum. Find the vertex. Find the x-intercepts. Describe the graph of the equation.

OpenStudy (anonymous):

\[y=x^2+6x-16\] like \[y=ax^2+bx+c\] open up because \(a>0\)

OpenStudy (anonymous):

I only need to Find the x-intercepts. AND Describe the graph of the equation.

OpenStudy (anonymous):

since it opens up vertex is a min

OpenStudy (anonymous):

vertex is \(-\frac{b}{2a}=-\frac{6}{2}=-3\) for the first coordinate, second one is -25

OpenStudy (anonymous):

x intercepts \[x^2+6x-16=0\] \[(x+8)(x-2)=0\] \[x=-8, x = 2\]

OpenStudy (anonymous):

to graph, put a parabola that opens up with vertex (-3,-25) and crosses x axis at (-8,0) and (2,0)